# Constant-roll inflation with tachyon field in the holographic braneworld

### v1

@@ 1 Introduction
The paper is organized as follows. In Section 2, we give a brief review of the tachyon dynamics in the holographic braneworld. The scenario of the constant-roll inflation in the holographic RSII model is discussed in Section 3. In Section 4, we present the details of our numerical results. In Section 5, we discuss the swampland criteria for the holographic constant-roll RSII inflation model. A summary and conclusions are given in Section 6.
2 The tachyon dynamics in the holographic braneworld
In this section, we briefly introduce the basic property of the tachyon dynamics in the holographic braneworld. In a braneworld scenario, our universe is a brane moving in the higher dimensional bulk, with only gravity allowed to propagate in the bulk. In the RSII model, two 3+1 dimensional branes with opposite tensions are embedded in a 4+1 spacetime with AdS5 geometry. The observers reside on the positive tension brane, and the negative tension brane is pushed off to infinity. If the brane is located at the boundary of an AdS5 space, the effective Einstein equations on the brane can be obtained in the framework of anti-de Sitter/conformal field theory (AdS/CFT) correspondence and holographic renormalization [18]. The effective Einstein equations yield the holographic Friedmann equations. For a spatially flat FLRW geometry on the brane, the holographic Friedmann equations are [14]
@@ 2 The tachyon dynamics in the holographic braneworld
where $p$ and $\rho$ are the pressure and the density of the fluid associated with the field on the brane and $h\equiv\ell H$ is a dimensionless Hubble expansion rate. The overdot in (2) denotes a derivative with respect to time measured in units of $\ell$. The fundamental coupling parameter $\kappa$ is related to the AdS5 curvature radius $\ell$ and the four-dimensional Newton constant $G_{\rm N}$
@@
In (2), we omit the contribution of ”dark radiation” [22]. The physically acceptable solution of the equation (1), as a quadratic equation for $h^{2}$, is of the form
$h^{2}=2\left(1-\sqrt{\frac{\kappa^{2}}{3}\ell^{4}\rho}\right).$ (4)
In the low-density limit ($\kappa^{2}\ell^{4}\rho\ll 1$), the equation (4) reduces to the standard Friedmann equation, ensures that the holographic cosmology reduces to the standard cosmology. According to (2), the expansion rate is a monotonously decreasing function of time. In this model, the evolution is constrained by the physical range of the expansion rate $0\leq h^{2}\leq 2$. The evolution of the universe starts with an initial $h_{\rm i}<\sqrt{2}$ and with negative $\dot{h}_{\rm i}$.
@@ 3 Constant-roll inflation
where $h_{*}$ is the expansion rate at an arbitrarily chosen time and $N$ is the number of e-folds defined as
@@
where $t_{\rm i}$ and $t_{\rm f}$ denote the times of the beginning and the end of inflation, respectively. To solve the most important problems in the early universe, the number of e-folds during inflation should be $N\simeq 60$. From the definition for the parameters $\varepsilon_{i}$ one obtains
@@
4 Numerical calculations
@@
In this section, we present the results for the observational parameters and the number of e-folds, obtained by numerically solving the equation for the Hubble expansion rate, equation (16), for fixed $\eta$. This equation is a second-order differential equation for the tachyon field $\theta$. To solve this equation numerically, we need to choose initial conditions, i.e., to fix the values of $h$ and $h_{,\theta}$ for a chosen value $\theta$. Let these be the values $h_{\rm i}\equiv h(\theta_{\rm i})$ and $h_{,\theta\rm i}\equiv h_{,\theta}(\theta_{\rm i})$ that correspond to the value of the field $\theta_{\rm i}$ at the time of the beginning of inflation ($\theta_{\rm i}=\theta(t_{\rm i})$). For this specific choice, we will fix the initial conditions from the values of the observational parameters. As we will see, this approach allows us to estimate the range of parameter $\eta$ in the model with the corresponding predictions that should be in good agreement with observational data. For numerical solving of (16), bearing in mind form of (14), without loss of generality, we can set $t_{\rm i}=0$ and $\theta_{\rm i}=0$.
@@
The observational parameters, the scalar spectral index $n_{\rm s}$, and the tensor-to-scalar ratio $r$ have been calculated in Ref. [24] for the tachyon model in the holographic braneworld up to the second order in the slow-roll parameters
@@
where $C\simeq-0.73$ and $\varepsilon_{i{\rm i}}\equiv\varepsilon_{i}(\theta_{\rm i})$ are the values of the slow-roll parameters at the beginning of inflation . The expressions (26) and (27) at linear order in the slow-roll parameters $\varepsilon_{i}$ agree with the expressions obtained in the standard scalar field inflation with a tachyon field (see e.g. ref. [23])
@@ 4.1 Numerical results
In Fig. 1 we plot $\eta$ and $h_{,\theta\rm i}$ as function of $h_{\rm i}$ for fixed and arbitrary $n_{\rm s}=0.966$ and $r=0.022$. As we shall see, for the values of the observational parameters, which are well inside the observational region, the model can provide the inflation stage with $N\approx 60$. We will use these values to illustrate further results.
@@
Now, we calculate the number of e-folds associated with the observational parameters, the values of which we choose according to the restrictions from the observational data. Using (14) the expression (20) becomes
[…]
where $\theta_{\rm f}=\theta(t_{\rm f})$ is fixed by $\varepsilon_{1}(\theta_{\rm f})=1$. To calculate the number of e-folds, we proceed as follows. For some chosen $n_{\rm s}$, $r$ and $h_{\rm\rm i}$, using (32) and (33), we calculate the parameter $\eta$ and the initial value $h_{,\theta\rm i}$. Then, we obtain $h(\theta)$ solving (15), for given $\eta$, with the initial conditions $h_{\rm i}$ and $h_{,\theta\rm i}$. From the condition for the end of inflation ($\varepsilon_{1}(\theta_{\rm f})=1$) we find $\theta_{\rm f}$. Finally, we use numerical integration in expression (34) to find $N$. For $n_{\rm s}=0.966$, $r=0.022$ and $h_{\rm i}=1.4$ the number of e-folds is $N=60.2$.
@@
4.2 Scalar spectral index and tensor-to-scalar ratio

### v2

@@ 1 Introduction
The paper is organized as follows. In Section 2, we give a brief review of the tachyon dynamics in the holographic braneworld. The scenario of the constant-roll inflation in the holographic model is discussed in Section 3. In Section 4, we briefly review the calculation of the primordial power spectra and present our numerical results. In Section 5, we discuss the swampland criteria for the holographic constant-roll RSII inflation model. A summary and conclusions are given in Section 6.
2 Tachyon cosmology in the holographic braneworld
We will consider two types of braneworlds in the AdS5 bulk: the second Randall-Sundrum (RSII) model and the holographic braneworld. In the RSII model [35], a 3-brane is located at a finite distance from the boundary of AdS5. The model was originally proposed as an alternative to the compactification of extra dimensions. In a holographic braneworld, a 3-brane is located at the boundary of the asymptotic AdS5 and the effective Einstein equations on the brane are derived in the framework of anti-de Sitter/conformal field theory (AdS/CFT) correspondence and holographic renormalization [36].
2.1 Randall-Sundrum and and holographic cosmologies
The first Randall-Sundrum model, usually referred to as the RSI model [34], is proposed as a solution to the hierarchy problem. The RSI model consists of two branes of opposite tensions embedded in the AdS5 bulk, where the observer brane is the brane with negative tension. In the RSII model, the observers are placed on a single positive tension brane. In a holographic braneworld universe, a 3-brane is located at the boundary of the asymptotic AdS5. In this way, the holographic cosmology may be regarded as a limiting case of the RSII cosmology. The braneworld univerrse evolution is governed by matter on the brane in addition to the CFT fluid.
A cosmology on the brane is obtained by allowing the brane to move in the bulk along the fifth dimension $z$. Equivalently, the brane is kept fixed at $z=z_{\rm br}$ while making the metric in the bulk time dependent. The time-dependent bulk spacetime with line element
$ds_{(5)}^{2}=\frac{\ell^{2}}{z^{2}}\left(n^{2}(\tau,z)d\tau^{2}-a^{2}(\tau,z)d\Omega_{\kappa}^{2}-dz^{2}\right),$ (1)
may be regarded as a $z$ foliation of the bulk with an FRW cosmology on each $z$ slice. In particular, at $z=z_{\rm br}$ we have the RSII cosmology and, at z=0, the holographic cosmology. The Friedmann equation on the RSII brane is modified
$H_{\rm br}^{2}+\frac{\kappa}{a_{\rm br}^{2}}=\frac{8\pi G_{\rm N}}{3}\rho_{\rm br}+\left(\frac{4\pi G_{\rm N}\ell}{3}\right)\rho_{\rm br}^{2}+\frac{\mu\ell^{2}}{a_{\rm br}^{4}},$ (2)
where $\ell$ is the AdS5 curvature radius, $H_{\rm br}=\dot{a}_{\rm br}/(n_{\rm br}a_{\rm br})$ is the Hubble rate on the RSII brane. Equation (2) was derived assuming the fine-tuned value of the brane tension $\lambda=3m_{\rm Pl}^{2}/(4\pi\ell^{2})$ which assures the vanishing of the cosmological constant. In the last term on the right-hand side of (2), usually referred to as the ”dark radiation”, the parameter $\mu$ is related to the bulk black-hole mass $\mu=(8G_{5}M_{\rm bh})/(3\pi\ell^{2}$ (see, e.g, [37]).
By introducing the boundary in AdS5 at $z=z_{\rm br}$ instead of $z=0$, the RSII model is conjectured to be dual to a cutoff CFT coupled to gravity, with $z=z_{\rm br}$ providing the IR cutoff. This conjecture then reduces to the standard AdS/CFT duality as the boundary is pushed off to $z=0$. In this limit, the variation of the action yields Einstein’s equations on the boundary [18, 19]
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=8\pi G_{\rm N}(\langle T^{\rm CFT}_{\mu\nu}\rangle+T^{\rm matt}_{\mu\nu}),$ (3)
where $T^{\rm matt}_{\mu\nu}$ is the energy-momentum tensor associated with matter on the holographic brane and $T^{\rm CFT}_{\mu\nu}$ the energy-momentum tensor of the CFT on the boundary. According to the AdS/CFT prescription, the expectation value $\langle T^{\rm CFT}_{\mu\nu}\rangle$ is obtained by functionally differentiating the renormalized on-shell bulk gravitational action with respect to the boundary metric [36]. From (3), one derives the Friedmann equation at the boundary
$\mathcal{H}^{2}=\frac{\ell^{2}}{4}\left(\mathcal{H}^{4}+\frac{4\mu}{a^{4}}\right)+\frac{8\pi G_{\rm N}}{3}\rho.$ (4)
where $\mathcal{H}^{2}=H^{2}+\kappa/a^{2}$ and $H=\dot{a}/a$ is the Hubble rate at the holographic boundary. As in the RSII model, the parameter $\mu$ is related to the black hole (BH) in the bulk, and the term proportional to $\mu$ is dubbed the ”dark radiation”. The second Friedmann equation may be obtained by combining the time derivative of (4) with the energy conservation equation
$\partial_{t}\rho+3(\rho+p)\frac{\partial_{t}a}{a}=0.$ (5)
One finds
$\frac{\ddot{a}}{a}\left(1-\frac{\ell^{2}}{2}\mathcal{H}^{2}\right)+\mathcal{H}^{2}=\frac{4\pi G_{\rm N}}{3}(\rho-3p).$ (6)
A map between RSII and holographic cosmologies can be constructed using [19]
$a_{\rm br}^{2}=a^{2}\left[\left(1-\frac{\mathcal{H}^{2}z^{2}}{4}\right)^{2}+\frac{1}{4}\frac{\mu z^{4}}{a^{4}}\right],\quad n_{\rm br}=\frac{\dot{a}_{\rm br}}{\dot{a}}.$ (7)
The Hubble rates are related by
$\mathcal{H_{\rm br}}\equiv H_{\rm br}^{2}+\frac{\kappa}{a_{\rm br}^{2}}=\mathcal{H}\frac{a}{a_{\rm br}}.$ (8)
Using this and (7) we can find a relation between the cosmological scales $a_{\rm br}$ on the brane at $z=z_{\rm br}$ and $a_{z}$ at on an arbitrary $z$-slice. First, we can express the first equation in (7) as an equation for $a^{2}$, $a_{z}^{2}$, and $\mathcal{H}_{z}^{2}$, and similarly as another equation for $a^{2}$, $a_{\rm br}$, and $\mathcal{H}_{\rm br}^{2}$ at $z=z_{\rm br}$. By eliminating $a^{2}$ from these two equations we find a relation between the scale $a_{z}$ and $a_{\rm br}$. Then, using this relation and (8) we can express the Hubble rate $H$ at $z=0$ in terms of the Hubble rate $H_{\rm br}$ at $z=z_{\rm br}$
$\mathcal{H}^{2}=2\mathcal{H}_{\rm br}^{2}\left(1+\frac{\mathcal{H}_{\rm br}^{2}z_{\rm br}^{2}}{2}\pm\sqrt{1+\mathcal{H}_{\rm br}^{2}z_{\rm br}^{2}-\frac{\mu z_{\rm br}^{4}}{a_{\rm br}^{4}}}\right)^{-1}.$ (9)
This equation provides a direct link between RSII and holographic cosmologies.
2.2 The tachyon dynamics in the holographic braneworld
In this section, we briefly introduce the basic property of the tachyon dynamics in the holographic braneworld. For a spatially flat FLRW geometry on the holographic brane, the modified Friedmann equations (4) and (6) can be written as
@@ 2.2 The tachyon dynamics in the holographic braneworld
where $p$ and $\rho$ are the pressure and the density of the fluid associated with the field on the brane and
$h\equiv\ell H.$ (12)
is a dimensionless Hubble expansion rate. The overdot in (11) denotes a derivative with respect to time measured in units of $\ell$. The fundamental coupling parameter $\kappa$ is related to the AdS5 curvature radius $\ell$ and the four-dimensional Newton constant $G_{\rm N}$
@@
As mentioned previously, Eqs. (10) and (11) might contain the term responsible for the so called ”dark radiation” and the cosmological constant term [18, 22]. We have omitted these terms for the following reasons. First, as in RSII cosmology, dark radiation is related to the BH in the AdS bulk. Since dark radiation is not relevant for inflation, we assume that the bulk is pure AdS5 with no BH, and hence, the terms responsible for it do not appear in our modified Friedmann equations (10) and (11). Second, by invoking the so-called RSII fine-tuning condition, we have adjusted the brane tension so that the cosmological constant vanishes. In principle, as discussed in Ref. [18], one could fine-tune the brane tension so that a tiny nonvanishing cosmological constant remains. This type of fine tuning may be relevant for recent cosmology. However, during inflation, the contribution of the cosmological constant is negligible since the tachyon field causes a fast quasi-de Sitter expansion anyway.
The physically acceptable solution to the equation (10), as a quadratic equation for $h^{2}$, is of the form
$h^{2}=2\left(1-\sqrt{1-\frac{\kappa^{2}}{3}\ell^{4}\rho}\right).$ (14)
In the low-density limit ($\kappa^{2}\ell^{4}\rho\ll 1$), this equation reduces to the standard Friedmann equation, to wit, the holographic cosmology reduces to the standard cosmology.
As discussed in Section 2.1 (for more details, see Ref. [18]), the holographic model is a limiting case of the RSII model with a single positive tension brane approaching the boundary of AdS5. Inflation based on the RSII model can have arbitrary large (up to $m_{\rm Pl}$) Hubble scale, depending on the ratio $\rho/\lambda$, where $\lambda$ is the brane tension (see, e.g., Ref. [37]). For exaample, if $\rho\sim\lambda$, one finds $H\sim 3/\ell$. In contrast, our Hubble scale is constrained by Eq. (14) and cannot exceed $H_{\rm max}\equiv\sqrt{2}/\ell$. According to our estimate (Section 4.4.1 further ahead), $\ell\sim 4\times 10^{5}l_{\rm Pl}$, and hence, the maximal Hubble scale in our holographic model $H_{\rm max}\sim 10^{-6}m_{\rm Pl}$ is way below the Planck scale.
According to (11), the expansion rate is a monotonously decreasing function of time. In this model, the evolution is constrained by the physical range of the expansion rate $0\leq h^{2}\leq 2$. The evolution of the universe starts with an initial $h_{\rm i}<\sqrt{2}$ and with negative $\dot{h}_{\rm i}$.
2.3 Tachyon fluid
@@ 3 Constant-roll inflation
where $h_{*}$ is the expansion rate at an arbitrarily chosen time and $N$ is the number of e-folds at time $t$ defined as
@@
where $t_{\rm i}$ denotes the time at the beginning of inflation. Using (24), Eq. (30) may be expressed as
[…]
To solve the most important problems in the early universe, the number of e-folds at the end of inflation should be $N\equiv N(\theta_{\rm f})\simeq 60$, where $\theta_{\rm f}=\theta(t_{\rm f})$ is fixed by $\varepsilon_{1}(\theta_{\rm f})=1$.
From the definition for the parameters $\varepsilon_{i}$ one obtains
@@
4 Primordial power spectra and observational parameters
Following Ref. [24], we first briefly review the calculations of the primordial power spectra for a general k-essence inflation in the holographic braneworld.
4.1 Scalar perturbations
Assuming a spatially flat background, we introduce the perturbed line element in the Newtonian gauge
$ds^{2}=(1+2\Psi)dt^{2}-(1-2\Phi)a^{2}(t)(dr^{2}+r^{2}d\Omega^{2}).$ (37)
Inserting the above metric components in the field equations, we obtain the Einstein equations at linear order. As shown in Ref. [24], one finds $\Phi/\Psi=1$ as in general relativity, and hence, one can work in longitudinal gauge with $\Psi=\Phi$. Then, the relevant Einstein equations at linear order are given by
$a^{-2}\nabla^{2}\Phi-3H\dot{\Phi}+3H^{2}\Phi=4\pi G_{\rm N}\delta{T}^{0}_{0}(1-h^{2}/2)^{-1},$ (38)
$(\dot{\Phi}+H\Phi)_{,i}=4\pi G_{\rm N}\delta{T}^{0}_{i}(1-h^{2}/2)^{-1}.$ (39)
The perturbations of the stress tensor components $\delta{T}^{\mu}_{\nu}$ are obtained by employing the perfect fluid description and the energy-momentum conservation. One finds [39]
$\delta{T}^{0}_{0}=\frac{{p}+{\rho}}{{c}_{\rm s}^{2}}\left[\left(\frac{\delta\theta}{\dot{\theta}}\right)^{.}-\Phi\right]-3H({p}+{\rho})\frac{\delta\theta}{\dot{\theta}}.$ (40)
$\delta{T}^{0}_{i}=({p}+{\rho})\left(\frac{\delta\theta}{\dot{\theta}}\right)_{,i}\,.$ (41)
As in Ref. [39], we introduce new functions
$\xi=\frac{a\Phi}{4\pi G_{\rm N}H},\quad\zeta=\Phi+H\frac{\delta\theta}{\dot{\theta}}.$ (42)
The quantity $\zeta$ is gauge invariant and measures the spatial curvature of comoving (or constant-$\theta$) hyper-surfaces. Then, using (40) and (41), equations (38) and (39) take the form
$a\dot{\xi}=z^{2}c_{\rm s}^{2}\zeta,$ (43)
$a\dot{\zeta}=z^{-2}\nabla^{2}\xi.$ (44)
Here
$z=\frac{a(\tilde{p}+\tilde{\rho})^{1/2}}{c_{\rm s}H}=\frac{a}{c_{\rm s}}\sqrt{\frac{\varepsilon_{1}}{4\pi G_{\rm N}}},$ (45)
where we have defined
$\tilde{p}+\tilde{\rho}=(p+\rho)(1-h^{2}/2)^{-1}.$ (46)
[…]
$v^{\prime\prime}-c_{\rm s}^{2}\nabla^{2}v-\frac{z^{\prime\prime}}{z}v=0,$ (47)
where the primes denote derivatives with respect to $\tau$. By making use of the Fourier transformation, we also obtain the mode-function equation
$v_{q}^{\prime\prime}+\left(c_{\rm s}^{2}q^{2}-\frac{z^{\prime\prime}}{z}\right)v_{q}=0.$ (48)
At linear order in $\varepsilon_{i}$ in the slow-roll regime, one can use the relations
$\tau=-\frac{1+\varepsilon_{1}}{aH}+\mathcal{O}(\varepsilon_{1}^{2}),\quad\frac{z^{\prime\prime}}{z}=\frac{\nu^{2}-1/4}{\tau^{2}},$ (49)
where
$\nu^{2}=\frac{9}{4}+3\varepsilon_{1}+\frac{3}{2}\varepsilon_{2}.$ (50)
We normalize the solution to (48) so that it satisfies the positive frequency asymptotic limit
$\lim_{\tau\rightarrow-\infty}v_{q}=\frac{e^{-ic_{\rm s}q\tau}}{\sqrt{2c_{\rm s}q}}..$ (51)
Then the solution which up to a phase agrees with (51) is given by
$v_{q}=\frac{\sqrt{\pi}}{2}(-\tau)^{1/2}H_{\nu}^{(1)}(-c_{\rm s}q\tau),$ (52)
where $H_{\nu}^{(1)}$ is the Hankel function of the first kind. This solution interpolates between subhorizon and superhorizon regimes, to wit, between short and long wavelength limits.
For a given wavenumber $q$, the potential term $z^{\prime\prime}/z$ in (48) can be neglected at sufficiently early times ($\tau\rightarrow-\infty$ ) when the physical wavelength of the perturbation $a/q$ is much smaller than the sound horizon $c_{\rm s}/H$. Hence, in this short wavelength limit, the normalized solutions are of the form (51).
The potential term $z^{\prime\prime}/z$ starts to dominate after these modes cross the sound horizon. Then, the solution to (48) evolves into the long wavelength solution
$v_{q}\simeq C_{q}z,$ (53)
where the constant $C_{q}$ is fixed by matching both solutions in the standard way. One finds
$|C_{q}|=\frac{1}{\bar{z}\sqrt{2\bar{c}_{\rm s}q}},$ (54)
where $\bar{z}$ and $\bar{c}_{\rm s}$ are the values of $z$ and $c_{\rm s}$ at the moment of sound horizon crossing $aH=c_{\rm s}q$. Hence, the curvature perturbation $\zeta_{q}=v_{q}/z\simeq C_{q}$ is conserved in the superhorizon region.
The calculation of the power spectrum proceeds as usual. Applying the standard canonical quantization [40], the field $v_{q}$ is promoted to an operator and the power spectrum of the field $\zeta_{q}$ is obtained from the two-point correlation function. The spectral density
$\mathcal{P}_{\rm S}(q)=\frac{q^{3}}{2\pi^{2}}|\zeta_{q}|^{2}=\frac{q^{3}}{2\pi^{2}z^{2}}|v_{q}|^{2},$ (55)
with $v_{q}$ given by (52), characterizes the primordial scalar fluctuations. Evaluating this at the horizon crossing one finds at the lowest order in $\varepsilon_{1}$ and $\varepsilon_{2}$
$\mathcal{P}_{{\rm S}}\simeq\frac{G_{\rm N}H^{2}}{\pi c_{\rm s}\varepsilon_{1}}\left[1-2\left(1+C\right)\varepsilon_{1}-C\varepsilon_{2}\right],$ (56)
where $C=\gamma-2+\ln 2\simeq-0.73$ and $\gamma$ is the Euler constant, so we recover the standard expression [23, 38].
4.2 Tensor perturbations
The tensor perturbations are related to the production of gravitational waves during inflation. The metric perturbation is, in this case, written as
$ds^{2}=dt^{2}-a^{2}(t)\left(\delta_{ij}+h_{ij}\right)dx^{i}dx^{j},$ (57)
where $h_{ij}$ is traceless and transverse. Inserting the metric components in the holographic field equations (see Ref. [24] for details) we obtain
$\left(1-\frac{\ell^{2}}{2}(H^{2}+\dot{H})\right)\left(\ddot{h}_{ij}+3H\dot{h}_{ij}-\frac{1}{a^{2}}\nabla^{2}h_{ij}\right)=8\pi G_{\rm N}\delta T_{ij}.$ (58)
In the absence of anisotropic stress, the gravitational waves decouple from matter, and the right-hand side of the above equation is zero. Then, the expression in the first bracket on the left-hand side factorizes out, and hence, our braneworld scenario does not introduce changes to the gravitational wave dynamics. At the lowest order in slow-roll parameters, the spectral density that characterizes the primordial tensor fluctuation is given by the usual expression [41]
$\mathcal{P}_{{\rm T}}\simeq\frac{16G_{\rm N}H^{2}}{\pi}\left[1-2\left(1+C\right)\varepsilon_{1}\right].$ (59)
4.3 Numerical calculations
@@
In this section, we present the results for the observational parameters and the number of e-folds, obtained by numerically solving the equation for the Hubble expansion rate, equation (26), for fixed $\eta$. This equation is a second-order differential equation with respect to the tachyon field $\theta$. To solve this equation numerically, we need to choose initial conditions, i.e., to fix the values of $h$ and $h_{,\theta}$ for a chosen value $\theta$. Let these be the values $h_{\rm i}\equiv h(\theta_{\rm i})$ and $h_{,\theta\rm i}\equiv h_{,\theta}(\theta_{\rm i})$ that correspond to the value of the field $\theta_{\rm i}$ at the time of the beginning of inflation ($\theta_{\rm i}=\theta(t_{\rm i})$). For this specific choice, we will fix the initial conditions from the values of the observational parameters. As we will see, this approach allows us to estimate the range of parameter $\eta$ in the model with the corresponding predictions that should be in good agreement with observational data. For numerical solving of (26), bearing in mind form of (24), without loss of generality, we can set $t_{\rm i}=0$ and $\theta_{\rm i}=0$.
@@
With the help of expressions (56) and (59), the observational parameters, the scalar spectral index $n_{\rm s}$, and the tensor-to-scalar ratio $r$ have been calculated in Ref. [24] for the tachyon model in the holographic braneworld. Up to the second order in the slow-roll parameters, one finds
@@
where $\varepsilon_{i{\rm i}}\equiv\varepsilon_{i}(\theta_{\rm i})$ are the values of the slow-roll parameters at the beginning of inflation. The expressions (60) and (61) at linear order in the slow-roll parameters $\varepsilon_{i}$ agree with the expressions obtained in the standard scalar field inflation with a tachyon field (see e.g. Ref. [23])
@@ 4.4 Numerical results
In Fig. 1 we plot $\eta$ and $h_{,\theta\rm i}$ as function of $h_{\rm i}$ for fixed and arbitrary $n_{\rm s}=0.966$ and $r=0.022$. As we shall see, for the values of the observational parameters, which are well inside the observational region, the model can provide the inflation stage with $N(\theta_{\rm f})\approx 60$. We will use these values to illustrate further results.
@@
Now, we calculate the number of e-folds associated with the observational parameters, the values of which we choose according to the restrictions from the observational data. Using (31) we define $N\equiv N(\theta_{\rm f})$, where $\theta_{\rm f}\equiv\theta(t_{\rm f})$ is fixed by $\varepsilon_{1}(\theta_{\rm f})=1$. To calculate the number of e-folds, we proceed as follows. For some chosen $n_{\rm s}$, $r$ and $h_{\rm\rm i}$, using (66) and (67), we calculate the parameter $\eta$ and the initial value $h_{,\theta\rm i}$. Then, we obtain $h(\theta)$ solving (26), for given $\eta$, with the initial conditions $h_{\rm i}$ and $h_{,\theta\rm i}$. From the condition for the end of inflation ($\varepsilon_{1}(\theta_{\rm f})=1$) we find $\theta_{\rm f}$. Finally, we use numerical integration in expression (31) to find $N$. For $n_{\rm s}=0.966$, $r=0.022$ and $h_{\rm i}=1.4$ the number of e-folds is $N=60.2$.
@@
4.4.1 Estimate of the scale
Here we estimate the radius of the AdS5 curvature $\ell$ by making use of the normalization of the curvature perturbation spectrum. We require that the curvature perturbation spectrum $\mathcal{P}_{\rm S}$ at the CMB pivot scale $q_{\rm CMB}=0.05{\rm Mpc}^{-1}$ coincides with the the power spectrum amplitude $A_{s}=2.10\pm 0.03\times 10^{-9}$ measured by Planck Collaboration 2018 [25]. For our purpose, we can assume $c_{\rm s}\simeq 1$ and approximate (56) by
$\mathcal{P}_{\rm S}(q)\simeq\frac{1}{\pi\varepsilon_{1}}\frac{H^{2}}{m_{\rm Pl}^{2}}=\frac{l_{\rm Pl}^{2}}{\pi\ell^{2}}\frac{h^{2}}{\varepsilon_{1}}$ (68)
where for each $q$, the quantities $h(\theta)$ and $\varepsilon_{1}(\theta)$ take on their horizon crossing values at the corresponding $N(\theta)$. In particular, we will use (68) at $q=q_{\rm CMB}$.
First, we find an approximate analytical function $N=N(\theta)$ using (31) and an approximate solution to (26) (deduced from Fig. 3)
$h\simeq h_{\rm i}e^{-\theta/(\alpha\ell)},$ (69)
where $\alpha$ is a constant. Then, from (31), we find
$N(\theta)\simeq\frac{3\alpha^{2}h_{\rm i}^{2}}{4}(1-e^{-2\theta/(\alpha\ell)}).$ (70)
Requiring $N(\theta_{\rm f}\simeq 60)$ at large $\theta_{\rm f}$ and taking $h_{\rm i}=1.4$, we find $\alpha=6.4$, and hence
$h\simeq 1.4e^{-\theta/(6.4\ell)},\quad N(\theta)\simeq 60(1-e^{-\theta/(3.2\ell)}).$ (71)
Next, for $q_{\rm CMB}$, we estimate the corresponding $N_{\rm CMB}\equiv N(\theta_{\rm CMB})$ at the horizon crossing, i.e., at the moment when
$a_{0}e^{N_{\rm CMB}}H_{\rm CMB}=q_{\rm CMB}$ (72)
where $H_{\rm CMB}\equiv H(\theta_{\rm CMB})$. We assume that the initial conditions at $a=a_{0}$ are set in the deep subhorizon region where $q_{\rm CMB}\gg a_{0}H_{\rm in}$. Hence, we set
$a_{0}H_{\rm in}=\epsilon q_{\rm CMB},$ (73)
where $\epsilon$ is a small parameter, e.g., of the order $\epsilon\sim 0.001$. Combining Eqs. (72) and (73) we obtain
$\epsilon e^{N_{\rm CMB}}\frac{H_{\rm CMB}}{H_{\rm in}}=1.$ (74)
Approximating $H_{\rm CMB}\simeq H_{\rm in}$, we obtain a rough estimate
$N_{\rm CMB}\simeq-\ln\epsilon\simeq 7.$ (75)
Then, from (71) we find $\theta_{\rm CMB}\simeq 0.4\ell$ and $h_{\rm CMB}\simeq 1.3$. Finally, using (68) and equating $\mathcal{P}(q_{\rm CMB})=2.1\times 10^{-9}$, we obtain an estimate
$\ell\simeq 4\times 10^{5}l_{\rm Pl}.$ (76)
4.5 Comparison with the Planck data